OpenStudy (bloomlocke367):

what's the standard equation of an ellipse in standard form and how do I make one with the vertex and co-vertex?

OpenStudy (bloomlocke367):

@iGreen

OpenStudy (igreen):

No idea.. @hartnn

OpenStudy (bloomlocke367):

@ganeshie8 @aaronq @Compassionate @TheSmartOne

OpenStudy (bloomlocke367):

that so confused me...

OpenStudy (bloomlocke367):

this is the problem I'm trying to solve: Write an equation of an ellipse in standard form with the center at the origin and with the given characteristics. vertex at (-3,0) and co-vertex at (0,2)

OpenStudy (bloomlocke367):

wait, is it: x^2/3+Y^2/2=1?..

OpenStudy (mathmate):

Almost, the standard equation for an ellipse centred at origin with semi-axes a,b is \(\large \frac{x^2}{a^2}+\frac{y^2}{b^2}=1\)

OpenStudy (bloomlocke367):

nonono.. it's x^2/9+y^2/4=1... because they're supposed to be squared..

OpenStudy (bloomlocke367):

lol. I just realized that..

OpenStudy (mathmate):

Good job! :)

OpenStudy (bloomlocke367):

so it is that, and not x^2/4+y^2/9=1, right?

OpenStudy (mathmate):

You need to square the semi-axis, so it's your later equation.

OpenStudy (bloomlocke367):

which one? is the 9 under the x^2 or the y^2? @mathmate

OpenStudy (mathmate):

*axes (plural) x^2/9+y^2/4=1 is correct. See the "a^2" and "b^2" in the standard equation? they correspond to (-3)^2 and 2^2) a corresponds to the x-intercept and b corresponds to the y-intercept.

OpenStudy (bloomlocke367):

okay, thanks :)

OpenStudy (mathmate):

You're welcome! :)